find the x- and y- intercepts and graph the equation plotting the intercepts.
5x-4y=-10 please show all of your work

Answer by nerdybill(7384)

5x-4y=-10
.
to find x-intercepts, set y=0 and solve for x:
5x-4(0)=-10
5x=-10
x = -2
so, x-intercept is at (-2,0)
.
to find y-intercepts, set x=0 and solve for y:
5(0)-4y=-10
-4y=-10
4y=10
y = 10/4
y = 5/2
y = 2.5
so, y-intercept is at (0, 2.5)